Ruby Company produces a chair that requires 4 yards of material per unit. The standard price of one yard of material is $12.20. During the month, 5,600 chairs were manufactured, using 22,600 yards at a cost of $11.47 per yard. Determine the following: Enter favorable variances as negative numbers. a. Price Variance b. Quantity Variance c. Cost Variance
Answer #1
a. Material Price Variance $             (16,498) favorable
b. Material Quantity Variance $                  2,440 unfavorable
c. Material Cost Variance $             (14,058) favorable
MATERIAL PRICE VARIANCE (MPV)
Actual Quantity (AQ) 22,600 yards a
Actual Price (AP) $11.47 b
Actual Cost $259,222.00 c = a × b
Standard Price (SP) $12.20 d
Actual Quantity (AQ) 22,600 yards e
Standard Cost of Actual Quantity $275,720.00 f = d × e
Material Price Variance $(16,498.00) c - f

Alternatively; MPV = AQ × (SP - AP) = 22,600 yards × ($12.20 - $11.47) = $16,498 (favorable)

MATERIAL QUANTITY VARIANCE (MQV)
Actual Quantity (AQ) 22,600 yards a
Standard Price (SP) $12.20 b
Standard Cost of Actual Quantity $275,720.00 c = a × b
Standard Quantity per chair 4 yards d
Actual production of chairs 5,600 chairs e
Total Standard Quantity (SQ) 22,400 yards f = d × e
Standard Price (SP) $12.20 g
Standard Cost of Standard Quantity for Actual production $273,280.00 h = f × g
Material Quantity Variance $2,440.00 c - h

Alternatively; MQV = SP × (SQ - AQ) = $12.20 × (22,400 yards - 22,600 yards) = $ (2,440) unfavorable (more yards of material used than the specified standards)

  • The change in above two alternative formula is due to rearrangement.
MATERIAL COST VARIANCE
Actual Quantity (AQ)                         22,600 yards a
Actual Price (AP) $11.47 b
Actual Cost $259,222.00 c = a × b
Standard Quantity per chair 4 yards d
Actual production of chairs 5,600 chairs e
Total Standard Quantity (SQ)                         22,400 yards f = d × e
Standard Price (SP) $12.20 g
Standard Cost $273,280.00 h = f × g
Material Cost Variance $             (14,058.00) c - h